OpenStudy (anonymous):

Rationalize the denominator and simplify. a) 12/√7 + 2 b) 6/√2 - √3 c) 15√6/3√5

OpenStudy (jdoe0001):

use the conjugate and multiply top and bottom by the conjugate of the denominator, for the first 2 the last one, just multiply by root 5 \(\bf \cfrac{12}{\sqrt{7}+2}\cdot \cfrac{\sqrt{7}-2}{\sqrt{7}-2}\implies \cfrac{12(\sqrt{7}-2)}{(\sqrt{7}+2)(\sqrt{7}-2)} \\ \quad \\ recall\implies {\color{blue}{ (a-b)(a+b) = a^2-b^2}}\qquad thus \\ \quad \\ \cfrac{12(\sqrt{7}-2)}{(\sqrt{7}+2)(\sqrt{7}-2)}\implies \cfrac{12(\sqrt{7}-2)}{(\sqrt{7})^2-2^2}\\ ---------------------------\\ \cfrac{6}{\sqrt{2}-\sqrt{3}}\cdot \cfrac{\sqrt{2}+\sqrt{3}}{\sqrt{2}+\sqrt{3}}\implies \cfrac{6(\sqrt{2}+\sqrt{3})}{(\sqrt{2}-\sqrt{3})(\sqrt{2}+\sqrt{3})}\implies \cfrac{6(\sqrt{2}+\sqrt{3})}{(\sqrt{2})^2-(\sqrt{3})^2}\\ ---------------------------\\ \cfrac{15\sqrt{6}}{3\sqrt{5}}\cdot \cfrac{3\sqrt{5}}{3\sqrt{5}}\implies \cfrac{(15\sqrt{6})(3\sqrt{5})}{3(\sqrt{5})^2}\) expand and simplify